# An Adaptive Method for Contextual Stochastic Multi-armed Bandits with Rewards Generated by a Linear Dynamical System

### v2

@@ V Bandit Strategy
The strategy the learner will use when choosing an action $a\in[k]$ is based on the theorems and lemmas mentioned above which can be broken down into two parts: Part 1: selecting parameter $s$ and Part 2: selecting action $a\in[k]$. This is comprehensively shown in Algorithm 1. Fig. 2 illustrates Part 1 and Part 2 of the algorithm.
@@
Note that in lines 19 and 20 of Algorithm 1 implies that the $u_{a}$ is set with the maximum $s_{a}$ out of all the actions. Not implementing lines 19 and 20 will cause the algorithm to commit to the action with the highest $s_{a}$ value instead of exploring all the actions.
Algorithm 1 Adaptive Recursive least-squares Exploration of System (ARES)
1:  Input: $\delta,\alpha\in(0,1)$, $s_{N},\lambda,B_{R}>0$
2:  $t\leftarrow 1$
3:  $\mathscr{S}=\{0,1,\dots,s_{N}\}$
4:  for $a\in[k]$ do
5:     $\mathscr{T}_{a}\leftarrow\{\}$
6:     for $s=\{0,1,\dots,s_{N}\}$ do
7:        $M_{a}\leftarrow 0$
8:        $N_{a}\leftarrow 0$
9:        $V_{a}\leftarrow\lambda I_{ms+1\times ms+1}$
10:        $\hat{G}_{a}^{t}\left(s\right)\leftarrow\mathbf{0}_{(ms+1)\times 1}$
11:        $\mathbf{e}_{a}^{t}\left(\delta,s\right)\leftarrow 1/\epsilon$ where $\epsilon$ small
12:        $\hat{J}_{a}^{t}\left(s\right)=0$
13:     end for
14:  end for
15:  for $t\in[n]$ do
16:     for $a\in[k]$ do
17:        $s_{a}\leftarrow\underset{s\in\{0,1,\dots,s_{N}\}}{\arg\min}\hat{J}_{a}^{t}\left(s\right)$
18:     end for
19:     $s\leftarrow\underset{a\in[k]}{\max}~{}s_{a}$
20:     $u_{a}\leftarrow\begin{cases}\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}\Theta_{t}}&\mbox{ if }t\geq s_{a}\\ \sqrt{2\frac{\log\left(1/\delta\right)}{M_{a}}}&\mbox{ if }t<s_{a}\end{cases}$
21:     $a^{*}\leftarrow\underset{a\in\{1,2,\dots,k\}}{\arg\max}\begin{cases}\hat{G}_{a}^{t}\left(s_{a}\right)^{\top}\Theta_{t}+u_{a}&\mbox{ if }t\geq s_{a}\\ \sum_{t\in\mathscr{T}_{a}}\frac{X_{t}}{M_{a}}+u_{a}&\mbox{ if }t<s_{a}\end{cases}$
22:     Sample $(X_{t},\theta_{t})$ from (1)
23:     $\mathscr{T}_{a^{*}}\leftarrow\mathscr{T}_{a^{*}}\cup\{t\}$
24:     $M_{a^{*}}\leftarrow M_{a^{*}}+1$ if $t<s_{a^{*}}$
25:     $N_{a^{*}}\leftarrow N_{a^{*}}+1$ if $t\geq s_{a^{*}}$
26:     for $s\in\{0,1,\dots,\min(t,s_{N})\}$ do
27:        $V_{a^{*}}^{t+1}\left(s\right)\leftarrow V_{a^{*}}^{t}\left(s\right)+\Theta_{\tau}\Theta_{\tau}^{\top}$
28:        $\hat{G}_{a^{*}}^{t+1}\left(s\right)^{\top}\leftarrow\hat{G}_{a^{*}}^{t}\left(s\right)^{\top}$ $+X_{\tau}\Theta_{\tau}^{\top}V_{a^{*}}^{t}\left(s\right)^{-1}$
29:        $\zeta_{a^{*}}^{t+1}\left(s\right)\leftarrow\alpha\zeta_{a^{*}}^{t}\left(s\right)+\left(1-\alpha\right)\left|X_{t}-\hat{G}_{a^{*}}^{t}\left(s\right)^{\top}\Theta_{t}\right|$
30:        $\hat{J}_{a^{*}}^{t}\left(s\right)\leftarrow\zeta_{a^{*}}^{t}+\mathbf{e}_{a^{*}}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a^{*}}^{t}\left(s\right)\Theta_{t}}$
31:        Update $\mathbf{e}_{a^{*}}^{t}(\delta,s)$ based on (80)
32:     end for
33:  end for
@@ VI Evaluation Study
The proposed Algorithm 1 (ARES) is compared with an Oracle (where the chosen action $a\in[k]$ is the action $a$ that has the maximum $\hat{X}_{t}$ from the Kalman filter (4)), UCB [2], Sliding Window UCB [3], Rexp3 [9], Random (for each round, the action is chosen uniformly random), and PIES [12]. PIES is similar to ARES where in the first $ks$ rounds the algorithm sequentially chooses each action $a\in[k]$ to identify the each vector $G_{a}\left(s\right)$. After the $ks$ rounds, PIES then chooses the action based on what it predicts to return the highest reward. In addition, PIES is preset with a parameter $s\in\{1,2,\dots,9,10\}$ which is the fixed window size of the algorithm. Each action has the same $s\in\{1,2,\dots,9,10\}$ value. The parameters used for ARES are in Table I.
@@
Parameter Value
$\delta$: 0.1
$\lambda$: 1
$\alpha$: 0.99
$s_{N}$: 10
$\nu$: 0.1
$B_{R}$: $\max_{a}c_{a}^{\top}Pc_{a}+\sigma_{\eta}^{2}$
$B_{G}$: $100$
$B_{c}$: $1$
@@
In Figure 3, the plots on the left are the median percentage of regret decreased when using ARES over the comparison algorithm. For the plots on the right, each line is each algorithm’s regret value normalized with respect to the Oracle’s regret. Note that we only plot the PIES’s normalized regret with parameters $s=1$ and $s=10$ since parameter $s=1$ is where PIES has the highest regret value while parameter $s=10$ is where PIES has the lowest regret value. It appears that both PIES and ARES perform better than any of the other comparison algorithms. For $\psi=1,2,3$, there exists a parameter $s$ value such that PIES does better than ARES. Each subtitle contains the set $\psi$ value, the maximum $c_{a}^{\top}Pc_{a}+\sigma_{\eta}^{2}$ value (where $P$ is the steady-state Kalman filter (5) error covariance matrix), and the spectral radius $\rho\left(\Psi-\Psi KC_{\theta}\right)$. It appears that as $\max_{a}c_{a}^{\top}Pc_{a}+\sigma_{\eta}^{2}$ and $\rho\left(\Psi-\Psi KC_{\theta}\right)$ increase, ARES starts to perform better than PIES for any preset parameter $s=1,\dots,10$ value. For $\psi=1,3$, there exists a parameter $s$ value such that PIES performs better than ARES, where ARES’s regret is at most $10$% more than than PIES regret. Finally, it appears that ARES and PIES normalized regrets appear to get closer to 1 (the Oracle’s regret value) as $\psi$ is closer to 1. This implies that as $\psi$ increases, the environment for PIES and ARES becomes harder, i.e. the model $G_{a}\left(s\right)$ is more difficult to identify.
@@
The parameter $s$ value in ARES is chosen such that it minimizes $\hat{J}_{a}^{t}\left(s\right)$ (91), where $\hat{J}_{a}^{t}\left(s\right)$ consists of two terms: $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$ and $\nu\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)\Theta_{t}}$. The term $\nu\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)\Theta_{t}}$ is similar to the perturbation input $u_{a}$; therefore, we only have to analyze the impact of $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$. Recall that PIES identifies $G_{a}\left(s\right)$ and uses the identified vector predicting the reward $X_{t}$ for each action $a\in[k]$ given a set parameter $s$ value, which is similar to ARES. Therefore, analyzing the PIES’s reward prediction error for each parameter $s$ value gives intuition on $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$ In Figure 5, the left plots are the average errors for each $a=1$ for each parameter $s=1,\dots,10$. For the right plots, these are the average errors of action $a=2$ for each parameter $s=1,\dots,10$. As shown, it appears that ARES on average chooses the parameter $s$ value that minimizes the prediction error $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$ the most for each action. Figure 5 also provides intuition on why ARES chooses a larger parameter $s$ value for one action versus another action. For example, in the third lowest plot of Figure 5, action $a=1$ error for parameters $s=8,9,10$ are similar. However, for action $a=2$, its error for parameter $s=10$ is the smallest value. Since we penalize for choosing high parameter $s$ values with $\nu\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)\Theta_{t}}$, ARES chooses the smallest parameter $s$ value action $a=1$ which is $s=8$, while ARES chooses parameter $s=10$ for action $a=2$ since it is the value that decreases the prediction error the most. Therefore, the takeaway from Figure 5 is if the prediction errors are approximately the same for a subset of parameter $s$ values, ARES will choose the minimum parameter $s$ value from the subset.
@@
VI-A How tight is ARES Regret Upper Bound?
To study the tightness of the ARES regret upper bound derived in Corollary 1, we plot a ratio of the derived bound (81) (in the numerator) and the average computed regret values of ARES (in the denominator). We set $\delta=0.1$. In the top plot of Figure 6, the ratio of regrets are shown for $\psi=4$ while the left plot of Figure 6 is the ratio of regrets for $\psi=1$. Notice that in the left plot as parameter $s$ increases from $s=1$ to $s=10$ the differences between the ratios increase. However, for the right plot, the differences between the ratios of regrets for parameters $s=1,3,10$ are smaller. Therefore, this follows from what was observed in the earlier figures where parameter $s$ has a larger impact when $\psi$ is close to $1$. In addition, we can observe that the regret bound becomes tighter as $\psi$ as parameter $s$ has a larger impact, verifying why for system (97) with parameter $\psi=1$ ARES chooses larger parameter $s$ values.
Figure 6: Each plot is a ratio of the theoretical regret bound (81) and the computed average regret from the ARES simulations.

### v3

@@ V Bandit Strategy
The strategy the learner will use when choosing an action $a\in[k]$ is based on the theorems and lemmas mentioned above which can be broken down into two parts: Part 1: selecting parameter $s$ and Part 2: selecting action $a\in[k]$. This is comprehensively shown in Algorithm ARES. Fig. 2 illustrates Part 1 and Part 2 of the algorithm.
@@
Note that in lines 19 and 20 of Algorithm ARES implies that the $u_{a}$ is set with the maximum $s_{a}$ out of all the actions. Not implementing lines 19 and 20 will cause the algorithm to commit to the action with the highest $s_{a}$ value instead of exploring all the actions.
Algorithm:
Adaptive Recursive least-squares Exploration of System (ARES)
Input: $\delta,\alpha\in(0,1)$, $s_{N},\lambda,B_{R}>0$
  1. Step 1:
    Initialize: $t\leftarrow 1$, $\mathscr{S}=\{0,1,\dots,s_{N}\}$
  2. Step 2:
    For each $a\in[k]$:
    1. (a)
      $\mathscr{T}_{a}\leftarrow\{\}$
    2. (b)
      For each $s\in\{0,1,\dots,s_{N}\}$:
      1. (i)
        $M_{a}\leftarrow 0$
      2. (ii)
        $N_{a}\leftarrow 0$
      3. (iii)
        $V_{a}\leftarrow\lambda I_{ms+1\times ms+1}$
      4. (iv)
        $\hat{G}_{a}^{t}(s)\leftarrow\mathbf{0}_{(ms+1)\times 1}$
      5. (v)
        $\mathbf{e}_{a}^{t}(\delta,s)\leftarrow 1/\epsilon$ (where $\epsilon$ is small)
      6. (vi)
        $\hat{J}_{a}^{t}(s)\leftarrow 0$
  3. Step 3:
    For each $t\in[n]$:
    1. (a)
      For each $a\in[k]$:
      1. (i)
        $s_{a}\leftarrow\arg\min_{s\in\{0,1,\dots,s_{N}\}}\hat{J}_{a}^{t}(s)$
    2. (b)
      $s\leftarrow\max_{a\in[k]}s_{a}$
    3. (c)
      $u_{a}$ calculation:
      1. (i)
        If $t\geq s_{a}$:
        $u_{a}\leftarrow\mathbf{e}_{a}^{t}(\delta,s)\sqrt{\Theta_{t}^{\top}V_{a}^{t}(s)^{-1}\Theta_{t}}$
      2. (ii)
        If $t<s_{a}$:
        $u_{a}\leftarrow\sqrt{2\frac{\log(1/\delta)}{M_{a}}}$
    4. (d)
      $a^{*}$ calculation:
      1. (i)
        If $t\geq s_{a}$:
        $a^{*}\leftarrow\underset{a\in[k]}{\arg\max}\hat{G}_{a}^{t}(s_{a})^{\top}\Theta_{t}+u_{a}$
      2. (ii)
        If $t<s_{a}$:
        $a^{*}\leftarrow\underset{a\in[k]}{\arg\max}\sum_{t\in\mathscr{T}_{a}}\frac{X_{t}}{M_{a}}+u_{a}$
    5. (e)
      Sample $(X_{t},\theta_{t})$ from Eq (1)
    6. (f)
      $\mathscr{T}_{a^{*}}\leftarrow\mathscr{T}_{a^{*}}\cup\{t\}$
    7. (g)
      Update counts:
      1. (i)
        If $t<s_{a^{*}}$:
        $M_{a^{*}}\leftarrow M_{a^{*}}+1$
      2. (ii)
        If $t\geq s_{a^{*}}$:
        $N_{a^{*}}\leftarrow N_{a^{*}}+1$
    8. (h)
      For each $s\in\{0,1,\dots,\min(t,s_{N})\}$:
      1. (i)
        $V_{a^{*}}^{t+1}(s)\leftarrow V_{a^{*}}^{t}(s)+\Theta_{\tau}\Theta_{\tau}^{\top}$
      2. (ii)
        $\hat{G}_{a^{*}}^{t+1}(s)^{\top}\leftarrow\hat{G}_{a^{*}}^{t}(s)^{\top}\\ +X_{\tau}\Theta_{\tau}^{\top}V_{a^{*}}^{t}(s)^{-1}$
      3. (iii)
        $\zeta_{a^{*}}^{t+1}(s)\leftarrow\alpha\zeta_{a^{*}}^{t}(s)\\ +(1-\alpha)\left|X_{t}-\hat{G}_{a^{*}}^{t}(s)^{\top}\Theta_{t}\right|$
      4. (iv)
        $\hat{J}_{a^{*}}^{t}(s)\leftarrow\zeta_{a^{*}}^{t}(s)\\ +\mathbf{e}_{a^{*}}^{t}(\delta,s)\sqrt{\Theta_{t}^{\top}V_{a^{*}}^{t}(s)\Theta_{t}}$
      5. (v)
        Update $\mathbf{e}_{a^{*}}^{t}(\delta,s)$ based on Eq (33)
@@ 3. Step 3:
V-A How tight is ARES Regret Upper Bound?
To study the tightness of the ARES regret upper bound derived in Corollary 1, we first prove the regret lower bound. The regret lower bound provides a metric of how difficult it is to consistently select the optimal action $a_{t}^{*}$. To acquire this bound, we first provide the following Oracle which selects actions $c_{a}\in\mathcal{A}$ based on the following optimization problem
$\begin{array}[]{cc}\underset{a\in[k]}{\arg\max}&\left\langle c_{a},\tilde{z}_{t}\right\rangle\\ \mbox{ s.t. }&\tilde{z}_{t+1}=\Gamma\tilde{z}_{t}+\Gamma\tilde{K}\left(Y_{t}-\tilde{C}\tilde{z}_{t}\right)\end{array}$ (38)
where $\tilde{K}$, $\tilde{C}$, and their associated terms are defined as
$\tilde{K}$ $\triangleq\tilde{P}\tilde{C}^{\top}\left(\tilde{C}\tilde{P}\tilde{C}^{\top}+\tilde{R}\right)^{-1}$
$\tilde{P}$ $=\Gamma\tilde{P}\Gamma^{\top}+Q$
$~~~~~-\Gamma\tilde{P}\tilde{C}^{\top}\left(\tilde{C}\tilde{P}\tilde{C}^{\top}+\tilde{R}\right)^{-1}\tilde{C}\tilde{P}\Gamma^{\top}$
$\tilde{C}$ $\triangleq\begin{pmatrix}c_{1}^{\top}&c_{2}^{\top}&\dots&c_{k}^{\top}&C_{\theta}^{\top}\end{pmatrix}^{\top}$
$\tilde{R}$ $\triangleq\mbox{blockdiag}\left(\sigma^{2},\dots,\sigma^{2},R\right)$
[…]
The Oracle is the method that selects actions $c_{a}\in\mathcal{A}$ using the context $\theta_{t}$ and rewards $X_{t}$ for each corresponding action $c_{a}\in\mathcal{A}$. In other words, this method is not viable with our assumptions as it observes more information than the problem formulation allows. Using the Oracle action selection method in (38), the following theorem proves a lower bound on its lowest obtainable regret. Proof for Theorem 4 is found in Appendix E.
Theorem 4.
Let there regret $R_{n}$ (3). The lower bound for regret $R_{n}$ is the following inequality
$R_{n}\geq n\sum_{i\in[k]}\sum_{j\in[k]}\sqrt{\frac{2\left(c_{a_{j}}-c_{a_{i}}\right)^{\top}Z\left(c_{a_{j}}-c_{a_{i}}\right)}{\mbox{tr}\left(\Psi_{i|j}\right)^{2k-2}\left|\tilde{\Sigma}_{i|j}\right|}}$ (39)
where $\tilde{\Sigma}_{i|j},\Psi_{i|j}$ are defined to be
$\tilde{\Sigma}_{i|j}$ $\triangleq A_{i}\tilde{Z}A_{i}^{\top}-A_{i}\tilde{Z}A_{j}^{\top}\left(A_{j}ZA_{j}^{\top}\right)^{-1}A_{j}\tilde{Z}A_{i}^{\top}$ (40)
$\Psi_{i|j}$ $\triangleq\begin{pmatrix}\Sigma_{i|j}^{-1}&\Sigma_{i|j}^{-1}\Pi_{i|j}\\ \Pi_{i|j}^{\top}\Sigma_{i|j}^{-1}&\Pi_{i|j}^{\top}\Sigma_{i|j}^{-1}\Pi_{i|j}\end{pmatrix}$ (41)
which are based on the following defined terms
$\tilde{Z}$ $\triangleq Z-\tilde{P}$ (42)
$\begin{pmatrix}A_{i}\left(z_{t}-e_{t|t-1}\right)\\ A_{j}z_{t}\end{pmatrix}$ $\sim\mathcal{N}\left(\mathbf{0},\Sigma_{i,j}\right)$
$\Sigma_{i,j}$ $\triangleq\begin{pmatrix}A_{i}\tilde{Z}A_{i}^{\top}&A_{i}\tilde{Z}A_{j}^{\top}\\ A_{j}\tilde{Z}A_{i}^{\top}&A_{j}ZA_{j}^{\top}\end{pmatrix}$ (43)
$A_{i}\left(z_{t}-e_{t|t-1}\right)\mid A_{j}z_{t}$ $\sim\mathcal{N}\left(\Pi_{i|j}z_{t},\tilde{\Sigma}_{i|j}\right)$
$A_{i}$ $\triangleq\begin{pmatrix}c_{a_{i}}-c_{a_{1}^{\prime}}&\dots&c_{a_{i}}-c_{a_{k-1}^{\prime}}\end{pmatrix}^{\top}$
$\Pi_{i|j}$ $\triangleq A_{i}\tilde{Z}A_{j}^{\top}\left(A_{j}ZA_{j}^{\top}\right)^{-1}A_{j},$
and $Z$ solves the Lyapunov equation $Z=\Gamma Z\Gamma^{\top}+Q$.
From Theorem 4, the Oracle cannot obtain a regret rate lower than linear. This follows from what was proven in Theorem 1 in [9] where the authors have proven that no bandit algorithm can obtain a regret lower than linear in environments that linearly change. Therefore, based on the asymptotic rate obtained in Corollary 1, having a superlinear regret rate of at most $n^{3/2}c^{s}\log\left(n\right)$ is appropriate, which can be lowered based on the design of window-size $s$.
@@ VI Evaluation Study
The proposed Algorithm ARES is compared with an Oracle (where the chosen action $a\in[k]$ is the action $a$ that has the maximum $\hat{X}_{t}$ from the Kalman filter (4)), UCB [2], Sliding Window UCB [3], Rexp3 [9], Random (for each round, the action is chosen uniformly random), and PIES [12]. PIES is similar to ARES where in the first $ks$ rounds the algorithm sequentially chooses each action $a\in[k]$ to identify the each vector $G_{a}\left(s\right)$. After the $ks$ rounds, PIES then chooses the action based on what it predicts to return the highest reward. In addition, PIES is preset with a parameter $s\in\{1,2,\dots,9,10\}$ which is the fixed window size of the algorithm. Each action has the same $s\in\{1,2,\dots,9,10\}$ value. The parameters used for ARES are in Table I.
@@
$\delta$ $\lambda$ $\alpha$ $s_{N}$ $\nu$ $B_{R}$ $B_{G}$ $B_{c}$
0.1 1 0.99 10 0.1 $\max_{a}c_{a}^{\top}Pc_{a}+\sigma_{\eta}^{2}$ 100 1
@@
In Figure 3, the plots on the left are the median percentage of regret decreased when using ARES over the comparison algorithm. For the plots on the right, each line is each algorithm’s regret value normalized with respect to the Oracle’s regret. Note that we only plot the PIES’s normalized regret with parameters $s=1$ and $s=10$ since parameter $s=1$ is where PIES has the highest regret value while parameter $s=10$ is where PIES has the lowest regret value. It appears that both PIES and ARES perform better than any of the other comparison algorithms. For $\psi=1,2,3$, there exists a parameter $s$ value such that PIES does better than ARES. Each subtitle contains the set $\psi$ value, the maximum $c_{a}^{\top}Pc_{a}+\sigma_{\eta}^{2}$ value (where $P$ is the steady-state Kalman filter (5) error covariance matrix), and the spectral radius $\rho\left(\Psi-\Psi KC_{\theta}\right)$. It appears that as $\max_{a}c_{a}^{\top}Pc_{a}+\sigma_{\eta}^{2}$ and $\rho\left(\Psi-\Psi KC_{\theta}\right)$ increase, ARES starts to perform better than PIES for any preset parameter $s=1,\dots,10$ value. For $\psi=1,3$, there exists a parameter $s$ value such that PIES performs better than ARES, where ARES’s regret is at most 1010% more than than PIES regret. Finally, it appears that ARES and PIES normalized regrets appear to get closer to 1 (the Oracle’s regret value) as $\psi$ is closer to 1. This implies that as $\psi$ increases, the environment for PIES and ARES becomes harder, i.e. the model $G_{a}\left(s\right)$ is more difficult to identify.
@@
The parameter $s$ value in ARES is chosen such that it minimizes $\hat{J}_{a}^{t}\left(s\right)$ (36), where $\hat{J}_{a}^{t}\left(s\right)$ consists of two terms: $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$ and $\nu\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)\Theta_{t}}$. The term $\nu\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)\Theta_{t}}$ is similar to the perturbation input $u_{a}$; therefore, we only have to analyze the impact of $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$. Recall that PIES identifies $G_{a}\left(s\right)$ and uses the identified vector predicting the reward $X_{t}$ for each action $a\in[k]$ given a set parameter $s$ value, which is similar to ARES. Therefore, analyzing the PIES’s reward prediction error for each parameter $s$ value gives intuition on $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$ . In Figure 5, the left plots are the average errors for each $a=1$ for each parameter $s=1,2,\dots,10$. For the right plots, these are the average errors of action $a=2$ for each parameter $s=1,2,\dots,10$. As shown, it appears that ARES on average chooses the parameter $s$ value that minimizes the prediction error $\left|X_{t}-\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}\right|$ the most for each action. Figure 5 also provides intuition on why ARES chooses a larger parameter $s$ value for one action versus another action. For example, in the third lowest plot of Figure 5, action $a=1$ error for parameters $s=8,9,10$ are similar. However, for action $a=2$, its error for parameter $s=10$ is the smallest value. Since we penalize for choosing high parameter $s$ values with $\nu\mathbf{e}_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)\Theta_{t}}$, ARES chooses the smallest parameter $s$ value action $a=1$ which is $s=8$, while ARES chooses parameter $s=10$ for action $a=2$ since it is the value that decreases the prediction error the most. Therefore, the takeaway from Figure 5 is if the prediction errors are approximately the same for a subset of parameter $s$ values, ARES will choose the minimum parameter $s$ value from the subset.